1. Keq = 8.63x10-4 for a reaction. Which of the following must be true?

∆Gº=0
∆Gº=1
∆Gº>0
∆Gº<0

2. What is the value of Keq if ∆Gº=-9.48 kJ (R=8.314 J/mol-K)?

6700
1.00
1.01
45.9

3. Calculate ∆Gº for the reaction, C3H8(g)+5O2(g)→3CO2(g)+4H2O(g), where ∆Gfº=-23.5 kJ/mol for CH4(g), -394 kJ/mol for CO2(g), and -229 kJ/mol for H2O(g):

600 kJ
-600 kJ
-2070 kJ
2070 kJ
Answer #1

1.

Keq= 8.63x10^-4

G0 = - RT lnKeq

Keq is a positive value

G0 getting the negative value

So, G0 is less than zero

G0<0

The answer is D

2)

G0 = - 9.48 KJ

R= 8.314 J/mol-k

T = 25C = 25+273=298K

G0 = - RT lnKeq

- 9.48x10^3 = - 8.314x273xlnKeq

lnKeq = 3.826

Keq= e^3.826

Keq= 45.88

Keq= 45.9

3)

C3H8(g) + 5O2(g) → 3CO2(g) + 4H2O(g

G0rxn = [ 3xG0f of CO2 + 4xG0f of H2O] - [ G0f of C3H8 + 5xG0f of O2]

G0rxn =[ 3 x( - 394) + 4x( -229)] - [ - 23.5 + 5 x 0.0]

G0rxn = - 2074.5 KJ

according to the given values the answer is

G0rxn = - 2074.5 KJ

(OR)

The answer is C